# Shadows of generalised Hayward spacetimes : in vacuum and with plasma

### v1

@@ Abstract
The Hayward regular BH solution attempted to resolve the curvature singularity issue by entering the domain of non-singular spacetimes. Recently, Dutta Roy and Kar (Phys. Rev. D 106, 044028) expanded this solution to encompass a broader range of spacetimes. These spacetimes are constructed based on the Damour-Solodukhin prescription, which involves introducing different metric parameters in the $g_{tt}$ and $g_{rr}$ components of the original Hayward line element, and are characterized by two parameters ($\sigma,\kappa$). This generalization gives rise to both known and novel regular/singular BHs as well as various types of wormhole spacetimes. In this work, we explore the spacetimes that emerge for different values of ($\sigma,\kappa$) from the generalized Hayward metric, particularly focusing on their shadows in vacuum and when surrounded by plasma. Intriguingly, we observe the presence of both photon and anti-photon spheres for certain regular spacetimes. Our study highlights the differences in the shadows of different types of regular spacetime compared to those of the singular BH derived from the generalized Hayward metric and also sheds light on the impact of plasma on the shadow radius.
I Introduction
The theory of general relativity (GR), with black holes (BH) as one of the solutions, is widely regarded as the most successful theory of gravity. Until recently, observational tests of GR were predominantly based on probing the weak field limit of gravity [1]. The era of testing GR in the regime of strong gravity and hence the existence of BHs started with the discovery of gravitational waves [2] by the LIGO-Virgo-KAGRA collaboration and simultaneously with the first detection of the shadow of the supermassive BH in M87 galaxy by the Event Horizon Telescope (EHT) collaboration [3]. Despite the success of GR, it is plagued with problems: the existence of curvature singularity, issues with quantization of gravity, are a few among them. To tackle the issue of curvature singularity in GR, a possible solution exists in literature, that of non-singular spacetimes. These regular classes of solutions can be broadly divided into spacetimes with and without horizon like the regular BHs and wormholes, respectively. However, upon a closer inspection of these solutions, one realizes that the issue of singularity is traded for other issues, such as energy condition violating matter for wormholes [4, 5] or mass-inflation instabilities in regular BHs [6, 7, 8, 9], provided that these solutions are viewed in the context of GR. While there exists multiple works which show the existence of wormholes in alternative theories of gravity without the need of ‘exotic’ matter [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22] and possible ways of taming the mass-inflation instability of regular BHs [23, 24], the issues still remains to be conclusively addressed.
@@ I Introduction
BHs are ubiquitous in physics, and at the same time they are the simplest objects in GR. For this reason they are often termed as the hydrogen atom of GR. The simplistic nature of the BHs, along with their extreme gravitational fields, make them the perfect testing bed to examine GR at strong field regimes. Towards this direction, the EHT collaboration has drawn much attention, and shifted the focus of the entire observational community with its first observations via the Very Long Baseline Interferometry (VLBI) of the supermassive BHs at the centers of the galaxies M87 [3] and Sgr $A^{*}$ [25]. Interestingly, it is to be noted that the EHT observations have achieved the required frequency to resolve the horizon-scale emission from the vicinity of the supermassive BHs at the center of the galaxies. As a result, the observed light from these extreme compact objects includes photons whose trajectories near the horizon of the BHs are deviated from their actual travel paths and describes nearly bound spherical orbits around the BH. This humongous achievement was made possible by a global network of radio telescopes working in parallel to create a virtual earth-sized observatory. By capturing and analyzing the radio waves emitted by the surrounding accretion disk, the EHT has once again provided a confirmation of the theoretical predictions of photon paths within the theory of GR.
@@
Although GR has achieved significant successes, it also faces severe challenges on the theoretical front. One of the most problematic areas in GR is the presence of the singularities at the core of the BH spacetimes, which raises serious question about the credibility of the theory. The issue that GR is ultraviolet (UV)-incomplete is related to this fact. It is therefore proposed that GR requires modifications in the regions where the spacetime curvature becomes high. Such modifications is necessary if one requires a theory of gravity which is UV complete. One potential approach to address the singularity issue is to consider specific types of matter distributions, which will lead to singularity-free BH solutions. Bardeen first proposed the metric for such a nonsingular BH geometry [26]. Later on, by coupling GR with nonlinear electrodynamics (NLED), a plethora of regular BH solutions with electric and magnetic charges were derived [27, 28, 29, 30, 31, 32, 33, 34]. Later on, Hayward proposed a static spherically symmetric BH solution in regular space-time [35]. This metric describes an isolated regular four dimensional spherically symmetric spacetime and contains a length scale $\ell$, apart from which, there is only one parameter: the mass of the BH. At large distances the metric takes the Schwarzschild form, while at the origin it is regular and has de Sitter form. The distinguishing feature of such a metric is its scaling behaviour, i.e. the scaling transformation of the coordinates, parameters and the metric preserves the form of the metric.
@@
The literature related to shadows of different extreme compact objects including BHs is very rich and because of the present interest in EHT collaboration’s result, there has been a flurry of works in this direction. Here, we briefly mention some of the important works towards the development of the subject. The shadow of the Schwarzschild black hole was first discussed in [36, 37]. Bardeen [38], again, was the first to describe the nature of the shadow cast by a rotating Kerr black hole, consequently it was generalised in the case of a Kerr-Newman BH in [39]. The shadow of the Kerr black hole as well as a Kerr naked singularity was discussed in [40] by introducing two new observables. Going beyond the standard GR solutions of Einstein’s equations, the study of shadows in BHs in other theories had been discussed extensively in the literature, viz. Einstein-Maxwell-Dilaton-Axion black hole [41], Kerr-Taub-NUT black hole [42], shadows of static spherically symmetric and rotating Kerr BHs in modified gravity (MOG) theory [43], rotating braneworld black hole [44], rotating non-Kerr type BHs [45], BHs in an expanding Universe [46], BHs in dynamical Chern-Simons modified gravity [47] to name a few. For more detailed on the methods and results in calculating shadows, we refer the readers to the existing reviews [48, 49, 50]. On the other hand, the study of shadows in regular BH backgrounds has gained considerable interests in recent years. There has been a lot of works done in this direction, see [51, 52, 53, 54, 55, 56, 57] for more details.
@@
In our work, we take a pragmatic approach and address the observational aspects of a class of regular spacetime solutions called the generalized Hayward metric recently developed in Dutta Roy and Kar [58], which describes wormholes, regular BH spacetimes along with the singular BH depending on the choice of the metric parameters. The generalized Hayward metric is a perfect testing ground for comparing the shadows of the regular spacetimes to that of the singular BHs. To our knowledge, the present study incorporating different classes of compact objects, eg. regular BHs, wormholes and singular BHs together, is new. Even though studies on each different topics including wormhole shadows (see [59] for details on wormhole shadows and for more references on the topic) have been performed individually, the present approach to study all these under the same umbrella is new. It is known that certain regular spacetimes possess an anti-photon ring along with the usual photon ring. We explore in detail the occurrence of photon/anti-photon rings in all classes of geometries originating from the generalized Hayward metric and probe the capability of the shadow observations to distinguish the regular spacetimes from the singular BH.
@@
It is known that the shadow of a compact object is associated with the photons scattered from the innermost circular null orbits and, hence, is a characteristic of the compact object. However, this holds true assuming the trajectory of these photons are unaffected throughout their journey from the point of emission to the typical asymptotic observer. In a realistic scenario, compact objects are always surrounded by an interstellar medium, which has a non-trivial effect on the photon trajectories. Hence it is natural to incorporate the effects of the environment in the computations of shadow. As a first step to address these issues, we analytically handle the computations of shadow in the presence of a plasma surrounding the compact object. Although our analysis corresponds to a simplified scenario which are different from a real astrophysical environment, these analytic calculations allow us to understand the explicit dependencies of the shadow profile on the metric and plasma parameters in contrast to numerical simulations.
@@
To this extent, we follow the techniques described in [60] and restrict to non-magnetized cold plasma consisting of two fluids, which are modeled by ions and electrons. The Hamiltonian for the photon trajectories in the presence of plasma in a curved spacetime background has been done at various levels of sophistication. For example, the rigorous derivation of Hamiltonian for magnetized plasma was achieved by Breur and Ehlers in [61, 62]. For a much simpler case of non-magnetized plasma fluids that we are interested in, the same analysis was carried out in [63, 64]. These results were then applied to the photon trajectors in the equatorial plane of Kerr spacetime by Perlick and others in [63, 65, 66, 67] and in a general setup in [68].
@@
The rest of the paper is organized as follows. In section II, we briefly discuss the generalized Hayward metric and the spacetimes arising for different metric parameter values. Section III summarizes the computation of the shadow cast by a static spherically symmetric solution. The main results are discussed in Section IV, consisting of the shadow, with and without plasma, for the different spacetimes originating from the generalized Hayward metric. We summarise the conclusions in Section V.
Unless otherwise stated we have used geometrised units $G=c=1$.
II Review of spacetimes originating from generalised Hayward metric
@@
$ds^{2}=-f_{1}(r)dt^{2}+\frac{dr^{2}}{f(r)}+r^{2}(d\theta^{2}+\sin^{2}\theta d\phi^{2}),$ (1)
@@
$f_{1}(r)=1-\frac{2M_{1}r^{2}}{r^{3}+2M_{1}\ell^{2}};\hskip 14.45377ptf(r)=1-\frac{2Mr^{2}}{r^{3}+2M\ell^{2}}.$ (2)
It is important to note that for $M_{1}=M$, this metric reduces to the Hayward regular BH, with the square of the parameter $\ell$ being identified with the inverse of the cosmological constant in the small $r$ limit. In fact, the construction of the generalized Hayward metric is motivated from the Hayward BH case by considering two different mass parameters in the $g_{tt}$ and $g_{rr}$ components. The incorporation of a different mass parameter $M_{1}$ in the $g_{tt}$ component of the metric, such that $M_{1}\neq M$, is inspired by the Damour-Solodukhin wormhole construction based on the Schwarzschild metric [69]. By introducing a different mass parameter into the metric of the Schwarzschild BH, the well-studied Damour-Solodukhin wormhole is obtained, while a procedure on the same vein on the Hayward regular BH resulted in the generalized Hayward metric as shown in eq.(1). Following [58], we re-write eq.(1) as,
$ds^{2}=-\Big{(}1-\frac{2\sigma x^{2}}{x^{3}+2\sigma\kappa^{2}}\Big{)}dt^{2}+\frac{dr^{2}}{\Big{(}1-\frac{2x^{2}}{x^{3}+2\kappa^{2}}\Big{)}}+r^{2}(d\theta^{2}+\sin^{2}\theta d\phi^{2}),$ (3)
where the parameters $M,M_{1},\ell$ are written in a dimensionless form using $\sigma=\frac{M_{1}}{M}$, $\kappa=\frac{\ell}{M}$ with the radial coordinate $x=\frac{r}{M}$. We will use the metric of the form given in eq.(3) for rest of the discussions. For different values of $(\sigma,\kappa)$, the spacetimes are classified into wormholes/BHs depending on the presence of throat/event horizon.
@@
Spacetime $\sigma$ $\kappa$ Singularity
@@
Table 1: Table shows the classification of spacetimes into BHs and wormholes depending on the values of ($\sigma,\kappa$). The only spacetime with singularity is the Schwarzschild BH while all others are regular.
III Null Geodesics and Shadow
This section mainly reproduces the known results and sets the conventions to be used in the later sections.In this section, we revisit the derivation of critical expressions for the shadow calculations. We will be broadly following the Hamilton-Jacobi approach to calculating the null geodesics. We restrict this analysis to the static spherically symmetric solutions with the ansatz being:
$ds^{2}=-f_{1}(r)dt^{2}+\frac{dr^{2}}{f(r)}+r^{2}d\Omega^{2},$ (4)
the choice of which is motivated by the form of the generalized Hayward metric, given in eq.(3), which is the focus of this work. The effective Lagrangian for the geodesic equations in the above spacetime is given by

### v2

@@ Abstract
We investigate the shadow properties of a wide class of spacetimes arising from different parameter regimes of the generalized Hayward metric, characterized by two independent parameters $(\sigma,\kappa)$ (Phys. Rev. D 106, 044028). This metric extends the original Hayward regular black hole solution by introducing distinct mass functions in the $g_{tt}$ and $g_{rr}$ components, giving rise to four types of wormholes ( which include multi-peak effective potentials), a regular black hole, and a singular black hole solutions allowing for a unified treatment of black hole mimickers. We compute the shadow radii for all spacetimes in vacuum and in the presence of plasma, using both homogeneous and non-homogeneous plasma profiles. Our results show that certain wormhole solutions particularly the Hayward-Damour-Solodukhin class—can exhibit multiple photon spheres, leading to shadow features that differ significantly from the Schwarzschild black hole. When these results are compared with Event Horizon Telescope observations of Sgr $A^{\star}$, we find that regular black holes remain observationally viable but only within a narrow parameter space. In contrast, wormhole solutions with multi-peak effective potentials are more consistent with shadow constraints than those with single peaks. This contrasts with quasinormal mode studies, which favored single-barrier potentials, and may imply detectable late-time echoes in gravitational wave signals.
1 Introduction
@@
Black holes (BH) are ubiquitous in physics, and at the same time they are the simplest objects in General Relativity(GR). For this reason they are often termed as the hydrogen atom of GR. Their extreme gravitational fields make them ideal candidates for testing GR in the strong-field regime. The onset of testing GR in the strong-field regmie began with the direct observations of gravitational waves by the LIGO-Virgo-KAGRA collaboration [1], which were closely followed by the Event Horizon Telescope’s (EHT) first observations via the Very Long Baseline Interferometry (VLBI) of the supermassive BHs at the centers of the galaxies M87 [2] and Sgr $\rm A^{*}$ [3]. It is to be noted that the EHT observations have achieved the required frequency to resolve the horizon-scale emission from the vicinity of the supermassive BHs at the center of the galaxies. As a result, the observed light from these extreme compact objects includes photons whose trajectories near the horizon of the BHs are deviated from their actual travel paths and describes nearly bound spherical orbits around the BH. This humongous achievement was made possible by a global network of radio telescopes working in parallel to create a virtual earth-sized observatory. By capturing and analyzing the radio waves emitted by the surrounding accretion disk, the EHT has once again provided a confirmation of the theoretical predictions of photon paths within the theory of GR.
Despite GR’s remarkable observational successes, it faces significant theoretical challenges, which include but are not limited to the complete understanding of curvature singularity and a comprehensive handle over the quantization of gravity. The issue that GR is ultraviolet (UV)-incomplete is related to this fact. It is therefore proposed that GR requires modifications in the regions where the spacetime curvature becomes high. Such modifications are necessary if one requires a theory of gravity which is UV complete.
@@
A potential approach to address the curvature singularity issue is to consider specific types of matter distributions, which will lead to singularity-free solutions. These regular solutions can be broadly divided into spacetimes with and without a horizon, like the regular BHs and traversable wormholes, respectively. Chronologically, Bardeen first proposed the metric for such a nonsingular BH geometry [4]. This was followed up by works on coupling GR with nonlinear electrodynamics (NLED) resulting in a plethora of regular BH solutions with electric and magnetic charges [5, 6, 7, 8, 9, 10, 11, 12]. One of the seminal works in the direction of obtaining a regular blackhole solution was by Hayward, who proposed a static spherically symmetric BH solution for which all possible curvature scalars are finite [13]. This metric describes an isolated regular four dimensional spherically symmetric spacetime and contains a length scale $\ell$, apart from which, there is only one parameter: the mass of the BH. At large distances the metric takes the Schwarzschild form, while at the origin it is regular and has de Sitter form. The distinguishing feature of such a metric is its scaling behaviour, i.e. the scaling transformation of the coordinates, parameters and the metric preserves the form of the metric.
However, a closer examination reveals that such ad hoc constructions often trade curvatures singularities for new issues. For instance, traversable wormholes typically require matter that violates energy conditions [14, 15], while regular black holes are prone to mass-inflation instabilities [16, 17, 18, 19] at least in the context of GR. However, beyond GR several modified gravity theories admit wormhole solutions without invoking exotic matter [20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32], and recent work has explored ways to mitigate the mass-inflation instability in regular black holes [33, 34]. Nevertheless, these issues remain far from conclusively resolved.
@@
The literature related to shadows of different extreme compact objects including BHs is very rich and because of the present interest in EHT collaboration’s result, there has been a flurry of works in this direction. Here, we briefly mention some of the important works towards the development of the subject. The shadow of the Schwarzschild black hole was first discussed in [35, 36]. Bardeen [37], again, was the first to describe the nature of the shadow cast by a rotating Kerr BH, consequently it was generalised in the case of a Kerr-Newman BH in [38]. The shadow of the Kerr BH as well as a Kerr naked singularity was discussed in [39] by introducing two new observables. Going beyond the standard GR solutions of Einstein’s equations, the study of shadows in BHs in other theories had been discussed extensively in the literature, viz. Einstein-Maxwell-Dilaton-Axion BH [40], Kerr-Taub-NUT BH [41], shadows of static spherically symmetric and rotating Kerr BHs in modified gravity (MOG) theory [42], rotating braneworld BH [43], rotating non-Kerr type BHs [44], BHs in an expanding Universe [45], BHs in dynamical Chern-Simons modified gravity [46] to name a few. For more detailed on the methods and results in calculating shadows, we refer the readers to the existing reviews [47, 48, 49]. On the other hand, the study of shadows in regular BH backgrounds has gained considerable interests in recent years. There has been a lot of works done in this direction, see [50, 51, 52, 53, 54, 55, 56] for more details.
@@
In our work, we take a pragmatic approach and address the observational aspects of a class of regular spacetime solutions called the generalized Hayward metric recently developed in Dutta Roy and Kar [57], which describes traversable wormholes, regular BH spacetimes along with the singular BH depending on the choice of the metric parameters. Note that the singularity or a lack thereof is defined with respect to curvature scalars. The generalized Hayward metric is a perfect testing ground for comparing the shadows of the regular spacetimes to that of the singular BHs.To our knowledge, this is the first study incorporating compact objects like regular BHs, wormholes, and singular BHs into a single framework. Even though studies on shadows of compact objects (see [58] for details on wormhole shadows and for more references on the topic) have been performed individually, the present approach to study a spacetime solution that changes its characteristics based on the parameter values has not been done before. As a part of this study, we also discovered that some of these cases possess an anti-photon ring along with the usual photon ring. We explore in detail the occurrence of photon/anti-photon rings in all classes of geometries originating from the generalized Hayward metric and probe the capability of the shadow observations to distinguish the regular spacetimes from the singular BH.
@@
It is known that the shadow of a compact object is associated with the photons scattered from the innermost unstable circular null orbits and, hence, is a characteristic of the compact object. However, this holds true assuming the trajectories of these photons are unaffected throughout their journey from the point of emission to the typical asymptotic observer. In a realistic scenario, compact objects are always surrounded by an interstellar medium, which has a non-trivial effect on the photon trajectories. Hence, it is natural to incorporate the effects of the environment in the computations of shadow. As a first step to address these issues, we analytically handle the computations of shadow in the presence of a plasma surrounding the compact object. Although our analysis corresponds to a simplified scenario which are different from a real astrophysical environment, these analytic calculations allow us to understand the explicit dependencies of the shadow profile on the metric and plasma parameters in contrast to numerical simulations.
@@
To this extent, we follow the techniques described in [59] and restrict to non-magnetized cold plasma consisting of two fluids, which are modeled by ions and electrons. The Hamiltonian for the photon trajectories in the presence of plasma in a curved spacetime background has been done at various levels of sophistication. For example, the rigorous derivation of Hamiltonian for magnetized plasma was achieved by Breur and Ehlers in [60, 61]. For a much simpler case of non-magnetized plasma fluids that we are interested in, the same analysis was carried out in [62, 63]. These results were then applied to the photon trajectories in the equatorial plane of Kerr spacetime by Perlick and others in [62, 64, 65, 66] and in a general setup in [67].
Finally, we address these predictions in the context of recent observational constraints from BH shadows of Sgr $A^{\star}$ from the EHT. Our results show that plasma effects, especially in non-homogeneous configurations, significantly constrain the viable parameter spaces of these regular spacetimes.
@@
The rest of the paper is organized as follows. In section 2, we briefly discuss the generalized Hayward metric and the spacetimes arising for different metric parameter values. Section 3 consists of a brief derivation of the shadow radius and other important quantities for the case of a static spherically symmetric solution in both vacuum and in the presence of a plasma medium. Section 4 consists of the shadow, with and without plasma, for the different spacetimes originating from the generalized Hayward metric. In section 5, we obtain additional constraints on the spacetime parameters and plasma parameter using EHT data for Sgr $A^{\star}$. We summarise the conclusions in Section 6. Unless otherwise stated we have used geometrised units $G=c=1$.
2 Review of spacetimes originating from generalised Hayward metric
@@
$ds^{2}=-f_{1}(r)dt^{2}+\frac{dr^{2}}{f(r)}+r^{2}(d\theta^{2}+\sin^{2}\theta d\phi^{2}),$ (2.1)
@@
$f_{1}(r)=1-\frac{2M_{1}r^{2}}{r^{3}+2M_{1}\ell^{2}};\hskip 14.45377ptf(r)=1-\frac{2Mr^{2}}{r^{3}+2M\ell^{2}}.$ (2.2)
It is important to note that for $M_{1}=M$, this metric reduces to the Hayward regular BH, with the square of the parameter $\ell$ being identified with the inverse of the cosmological constant in the small $r$ limit. In fact, the construction of the generalized Hayward metric is motivated from the Hayward BH case by considering two different mass parameters in the $g_{tt}$ and $g_{rr}$ components. The incorporation of a different mass parameter $M_{1}$ in the $g_{tt}$ component of the metric, such that $M_{1}\neq M$, is inspired by the Damour-Solodukhin wormhole construction based on the Schwarzschild metric [68]. By introducing a different mass parameter into the metric of the Schwarzschild BH, the well-studied Damour-Solodukhin wormhole is obtained, while a procedure on the same vein on the Hayward regular BH resulted in the generalized Hayward metric as shown in eq.(2.1). Following [57], we re-write eq.(2.1) as,
$ds^{2}=-\Big{(}1-\frac{2\sigma x^{2}}{x^{3}+2\sigma\kappa^{2}}\Big{)}dt^{2}+\frac{dr^{2}}{\Big{(}1-\frac{2x^{2}}{x^{3}+2\kappa^{2}}\Big{)}}+r^{2}(d\theta^{2}+\sin^{2}\theta d\phi^{2}),$ (2.3)
where the parameters $M,M_{1},\ell$ are written in a dimensionless form using $\sigma=\frac{M_{1}}{M}$, $\kappa=\frac{\ell}{M}$ with the radial coordinate $x=\frac{r}{M}$. We will use the metric of the form given in eq.(2.3) for rest of the discussions. For different values of $(\sigma,\kappa)$, the spacetimes are classified into wormholes/BHs depending on the presence of throat/event horizon.
@@
Spacetime $\sigma$ $\kappa$ Curvature Singularity
@@
Table 1: Table shows the classification of spacetimes into BHs and wormholes depending on the values of ($\sigma,\kappa$). The only spacetime with curvature singularity is the Schwarzschild BH while all others are regular.
3 Null Geodesics and Shadow
The ‘shadow’ in the context of BHs and other astrophysical compact objects is defined as a dark silhouette against a bright background which is formed due to the gravitational bending of light. Historically this has been called by various names and has various definition. A comprehensive list of these have been discussed in [48]. In the context of this paper we define shadow as the two dimensional closed curve on a celestial sphere that separates the captured and scattered orbits [69, 70, 36]. These celestial coordinates can be expressed in terms of the observer’s position relative to the BH and the tangents to the photon geodesics [47, 70, 71, 38]
$\alpha$ $=-r_{o}^{2}\sin\theta_{o}\left.\frac{d\phi}{dr}\right\rvert_{r_{o}},$ (3.1)
$\beta$ $=\left.r_{o}^{2}\frac{d\theta}{dr}\right\rvert_{r_{o}},$
where $\alpha$ and $\beta$ are celestial coordinates. In case of static spherically symmetric metric the closed curve is a circle and the radius of this circle is known as the shadow radius and is determined by
$r_{sh}^{2}=\alpha^{2}+\beta^{2}.$ (3.2)
Hence calculating shadow radius essentially reduces to finding the tangents to the null geodesics and plugging that information into the above equations. In principle there are various approaches for this calculation. In the remainder of this section we reproduce the known results and set the conventions to be used in the later sections of this work based on the Hamilton-Jacobi approach. In the process, we also define other useful quantities, such as the impact parameter and the photon sphere radius. We also restrict this analysis to the static spherically symmetric solutions with the metric ansatz being
$ds^{2}=-f_{1}(r)dt^{2}+\frac{dr^{2}}{f(r)}+r^{2}d\Omega^{2},$ (3.3)
the choice of which is motivated by the form of the generalized Hayward metric, given in eq.(2.3), which is the focus of this work. The effective Lagrangian for the geodesic equations in the above spacetime is given by